# Fast and Robust Stationary Crowd Counting with Commodity WiFi

### v1

@@ Abstract
This paper introduces a novel method for estimating the size of seated crowds with commodity WiFi signals, by leveraging natural body fidgeting behaviors as a passive sensing cue. Departing from prior binary fidget representations, our approach leverages the bandwidth of the received signal as a finer-grained and robust indicator of crowd counts. More specifically, we propose a mathematical model that relates the probability density function (PDF) of the signal bandwidth to the crowd size, using a principled derivation based on the PDF of an individual’s fidget-induced bandwidth. To characterize the individual fidgeting PDF, we use publicly available online videos, each of a seated individual, from which we extract body motion profiles using vision techniques, followed by a speed-to-bandwidth conversion inspired by Carson’s Rule from analog FM radio design. Finally, to enhance robustness in real-world deployments where unrelated motions may occur nearby, we further introduce an anomaly detection module that filters out non-fidget movements. We validate our system through 42 experiments across two indoor environments with crowd sizes up to and including 13 people, achieving a mean absolute error of 1.04 and a normalized mean square error of 0.15, with an average convergence time of 51 seconds, significantly reducing the convergence time as compared to the state of the art. Additional simulation results demonstrate scalability to larger crowd sizes. Overall, our results show that our pipeline enables fast, robust, and highly accurate counting of seated crowds.
@@ 1 Introduction
Passive human sensing and activity-aware smart environments have emerged as prominent areas of research in recent years. The widespread availability of sensing technologies (such as computer vision, radio frequency (RF), and acoustic) in everyday environments has enabled new opportunities for the ubiquitous monitoring of human presence and activities. The common goal of passive human sensing research is to enhance safety, healthcare, and quality of life without requiring direct interaction from the users. Researchers have utilized these modalities for a variety of applications, such as fall detection [1, 2, 3], early diagnosis of medical conditions and smart healthcare [4, 5, 6, 7], activity recognition [8, 9, 10], and occupancy estimation [11, 12, 13, 14], among others.
@@
To enable occupancy analytics, different sensing modalities have been used in recent years. Vision-based methods, for instance, have been utilized heavily for occupancy estimation tasks, particularly through the use of machine learning techniques. Common approaches include leveraging convolutional neural networks (CNNs) [15] or density maps [16] to perform crowd counting. To further enhance coverage and robustness, other studies have explored multi-camera systems that address occlusions and increase spatial visibility [17, 18]. Despite their success in crowd counting performance, vision-based systems face significant limitations, including privacy concerns, dependence on line-of-sight, and the need for consistent lighting, underscoring the need for alternative crowd sensing modalities.
@@
More recently, researchers have used radio frequency-based solutions for occupancy estimation and crowd counting to address the limitations of vision-based methods. For instance, [19] proposed a mmWave-based feature extraction crowd counting method, using a k-nearest-neighbor classifier. Similarly, [20] introduced mmCount, a stationary crowd-counting system based on 2D heatmaps generated with mmWave radar. Sakhnini et al. [21] proposed a mmWave-based metric correlated with large crowd sizes. Deep learning [22] and transformer-based [23] approaches have also gained attention for radar-based crowd counting. In general, although radar-based methods are suitable for specific applications, they often require either customized hardware or more expensive commodity devices as compared to WiFi. Moreover, they can experience severe blockage as compared to WiFi, limiting their practicality as a ubiquitous solution.
@@
In contrast, WiFi-based sensing has gained traction in the research community due to the widespread availability of WiFi infrastructure in everyday environments and its through-wall sensing capabilities. Earlier attempts in crowd counting  [24, 25] proposed a mathematical framework for counting moving crowds. More recent approaches [26, 27, 28, 29, 30, 31] have also used machine learning for this purpose, albeit with limited generalizability, as collecting diverse crowd training data is challenging. Overall, the aforementioned WiFi-based crowd counting work are focused on counting moving crowds, and the methods are not applicable to the problem of interest to this paper, which is focused on counting a seated crowd.
@@
In this paper, we are interested in counting seated crowds with WiFi signals, a relatively underexplored problem in the RF sensing literature, as it is considerably more challenging due to the lack of major body motion. In [32, 33, 34], neural network pipelines are trained to count a number of seated or mobile people. However, these work involve very small group sizes (e.g., up to five individuals) and are thus better categorized as multi-person counting rather than true crowd size estimation. They can also suffer from limited generalizability, as collecting diverse crowd training data is challenging. In recent years, stationary crowd counting with WiFi CSI has also been tackled using fidget duty cycles, incorporating queuing theory to model fidgeting statistics. Specifically, in [11], we proposed this approach to estimate the size of a seated crowd, while [35] adopts a similar methodology in the context of passenger counting. Using fidget-induced signal variations is a promising first step toward solving the challenging problem of counting seated crowds, where large body motions are absent. However, the aforementioned existing method typically requires long observation windows to accumulate sufficient fidgeting statistics, relying on the observation of several fidget and silent periods to form a stable estimate. Therefore, these binary duty cycle-based approaches, where signal variations are treated as on-off fidget activity indicators, lead to coarse-grained modeling that prolongs estimation convergence.
@@
In this paper, we take a different perspective for counting a seated crowd. More specifically, we propose a finer-grained approach that efficiently models the full amplitude of fidget-induced signal fluctuations, as opposed to an on-off duty cycle-based approach. This approach allows our system to capture more information about the fidgeting behavior of crowds and converge with significantly less observation time.
@@ 2 Problem Formulation
[…]
@@
[…]
@@
Figure 2: Sample WiFi experiment from a stationary crowd to demonstrate the difference between binary and non-binary bandwidth-based motion representations – (a) spectrogram of the received WiFi signal capturing real-time fidget activities, (b) binary fidget representation obtained via thresholding, which provides a coarse, on/off encoding of motion, (c) proposed bandwidth-based representation enabling a more fine-grained and informative motion description, if it could be related to the crowd size in a principled way. In this paper, we show how to methodically extract the crowd size from the bandwidth.
@@ 4.1 Relating Body Part Speeds of an Individual to Bandwidth
4.2 Modeling the Observed Bandwidth of a Stationary Crowd of Size $N$
@@
[…]
@@
$=\prod_{i=1}^{N}P\left(BW_{i}\leq x\right)\quad\text{(}BW_{i}\text{ are independent)}$)
$=\left(P\left(BW\leq x\right)\right)^{N}\mkern 1.0mu\text{(}BW_{i}\text{ are identically dist.)}$.)
@@ 5.1 Experiment Settings
We conducted our experiments in two distinct indoor environments: a conference room and an office space, as shown in Fig. 1. We chose indoor spaces, as they present considerably more challenges due to the higher level of multipath, as compared to outdoor spaces. Moreover, the two indoor spaces are further chosen to represent different degrees of multipath. The conference room is moderately furnished and exhibits a smaller level of multipath propagation, thus offering a relatively clean RF environment. In contrast, the office is densely populated with metallic furniture, shelves, and electronics, resulting in a highly cluttered environment with significant multipath effects. This diversity allows us to evaluate the robustness of our approach in both relatively challenging and considerably challenging RF conditions.
In each environment, we conducted a series of experiments with varying crowd sizes and seating arrangements. Specifically, a subset of 13 participants is recruited and asked to be seated for reading. The configuration of the seating, including the number of rows of chairs, spacing between individuals, and alignment with respect to the TX-RX link, is varied across trials to ensure generalization across spatial layouts. Each experiment lasted around 3 minutes. We note that the small variation around experiment times is due to experimental conditions, external interventions, and participant preparations.
@@ 6.2 Simulation-Based Results to Test Larger Crowd Sizes
So far, we have shown real experimental results with crowds up to and including 13 individuals. To showcase our pipeline’s performance with even larger crowds, we next present a novel framework to efficiently simulate and test with larger crowd sizes. Our pipeline efficiently simulates crowd bandwidths from readily-available videos. By using our video-to-bandwidth modeling, we can eliminate the need for a whole-body crowd wave simulator, which can be considerably challenging and computationally prohibitive. More specifically, to simulate synthetic bandwidth observations for larger crowd sizes, we use the video-to-bandwidth pipeline of Section 4.3 and public videos of single seated individuals. Each video is segmented into 3-minute non-overlapping segments. Each segment is then processed through the video-to-bandwidth pipeline, which yields the corresponding bandwidth measurements for the seated individual as if a transceiver were present. To synthetically generate test data for any crowd size $N$, we then randomly sample $N$ of these segments and combine them by taking the maximum bandwidth value across all $N$ segments at each time instance. This approach efficiently models the observed bandwidth expected from a group of $N$ independently fidgeting, seated individuals, where the overall observed bandwidth is dictated by the highest contributor at any given time. We repeat this sampling strategy for $N\in{2,3,...,20}$ with as many synthetic test samples generated per $N$ as the number of unique combinations that can be created from the video pool. Fig. 4 summarizes our proposed method for synthetic test data generation.
@@
In order to mitigate the bias towards any specific random selection, we repeat the entire cross-validation process five times with different randomizations and report the average results. In total, for each crowd size, we evaluate our pipeline on $296$ unique 3-minute segments per crowd size, yielding a total test set size of $20\times 296=5920$ empirical test instances. We note that each data segment is unique. However, during the extrapolation to larger crowd sizes, different segments belonging to the same video may be randomly picked as part of the same synthetic crowd data sample. Table 4 reports the MAE for each crowd size. Our system achieves an overall NMSE of $0.054$ and an overall MAE of $1.47$ over the simulated dataset. These results further showcase the potential of our bandwidth-based approach with even larger crowd sizes.

### v2

@@ Abstract
This paper introduces a novel method for estimating the size of seated crowds with commodity WiFi signals, by leveraging natural body fidgeting behaviors as a passive sensing cue. Departing from prior binary fidget representations, our approach leverages the bandwidth of the received signal as a finer-grained and robust indicator of crowd counts. More specifically, we propose a mathematical model that relates the probability density function (PDF) of the signal bandwidth to the crowd size, using a principled derivation based on the PDF of an individual’s fidget-induced bandwidth. To characterize the individual fidgeting PDF, we use publicly available online videos, each of a seated individual, from which we extract body motion profiles using vision techniques, followed by a speed-to-bandwidth conversion inspired by Carson’s Rule from analog FM radio design. Finally, to enhance robustness in real-world deployments where unrelated motions may occur nearby, we further introduce an anomaly detection module that filters out non-fidget movements. We validate our system through 42 experiments across two indoor environments with crowd sizes up to and including 13 people, achieving a mean absolute error of 1.04 and a normalized mean square error of 0.15, with an average convergence time of 51 seconds, significantly reducing the convergence time as compared to the state of the art. Additional simulation results demonstrate scalability to larger crowd sizes. Overall, our results showcase that the proposed pipeline enables fast, robust, and highly accurate counting of seated crowds.
@@ 1 Introduction
Passive human sensing and activity-aware smart environments have emerged as prominent areas of research in recent years. The widespread availability of sensing technologies (such as computer vision, radio frequency (RF), and acoustic) in everyday environments has enabled new opportunities for the ubiquitous monitoring of human presence and activities. The common goal of passive human sensing research is to enhance safety, healthcare, and quality of life without requiring direct interaction from the users. Researchers have utilized these modalities for a variety of applications, such as fall detection [1, 2, 3], early diagnosis of medical conditions and smart healthcare [4, 5, 6, 7], activity recognition [8, 9, 10], person identification [11, 12], and occupancy estimation [13, 14, 15, 16], among others.
@@
To enable occupancy analytics, different sensing modalities have been used in recent years. Vision-based methods, for instance, have been utilized heavily for occupancy estimation tasks, particularly through the use of machine learning techniques. Common approaches include using convolutional neural networks (CNNs) [17] or density maps [18] to perform crowd counting. To further enhance coverage and robustness, other studies have explored multi-camera systems that address occlusions and increase spatial visibility [19, 20]. Despite their success in crowd counting performance, vision-based systems face significant limitations, including privacy concerns, dependence on line-of-sight, and the need for consistent lighting, highlighting the need for alternative crowd sensing modalities.
@@
More recently, researchers have used radio frequency-based solutions for occupancy estimation and crowd counting to address the limitations of vision-based methods. For instance, [21] proposed a mmWave-based feature extraction method for crowd counting, using a k-nearest-neighbor classifier. Similarly, [22] introduced mmCount, a stationary crowd-counting system based on 2D heatmaps generated with mmWave radar. Sakhnini et al. [23] proposed a mmWave-based metric correlated with large crowd sizes. Deep learning [24] and transformer-based [25] approaches have also gained attention for radar-based crowd counting. In general, although radar-based methods are suitable for specific applications, they often require either customized hardware or more expensive commodity devices as compared to WiFi. Moreover, they can experience severe blockage as compared to WiFi, limiting their practicality as a ubiquitous solution.
@@
In contrast, WiFi-based sensing has gained traction in the research community due to the widespread availability of WiFi infrastructure in everyday environments and its through-wall sensing capabilities. Earlier attempts in crowd counting  [26, 27] proposed a mathematical framework for counting moving crowds. More recent approaches [28, 29, 30, 31, 32, 33] have also used machine learning for this purpose, albeit with limited generalizability, as collecting diverse crowd training data is challenging. Overall, the aforementioned WiFi-based crowd counting work are focused on counting moving crowds, and the methods are not applicable to the problem of interest in this paper, which is focused on counting a seated crowd.
@@
In this paper, we are interested in counting seated crowds with WiFi signals, a relatively underexplored problem in the RF sensing literature, as it is considerably more challenging due to the lack of major body motion. In [34, 35, 36, 37], neural network pipelines are trained to count a number of seated or mobile people. These work, however, can suffer from limited generalizability, as collecting diverse crowd RF training data is challenging. Moreover, most of the aforementioned work (except [37]) involve very small group sizes (e.g., up to five individuals) and are thus better categorized as multi-person counting rather than true crowd size estimation. In recent years, stationary crowd counting with WiFi CSI has also been tackled using fidget duty cycles, incorporating queuing theory to model fidgeting statistics. Specifically, in [13], we proposed this approach to estimate the size of a seated crowd, while [38] adopts a similar methodology in the context of passenger counting. Using fidget-induced signal variations is a promising first step toward solving the challenging problem of counting seated crowds, where large body motions are absent. However, the aforementioned existing method typically requires long observation windows to accumulate sufficient fidgeting statistics, relying on the observation of several fidget and silent periods to form a stable estimate. Therefore, these binary duty cycle-based approaches, where signal variations are treated as on-off fidget activity indicators, lead to coarse-grained modeling that prolongs estimation convergence.
@@
In this paper, we take a different perspective on counting a seated crowd. More specifically, we propose a finer-grained approach that efficiently models the full amplitude of fidget-induced signal fluctuations, as opposed to an on-off duty cycle-based approach. This approach allows our system to capture more information about the fidgeting behavior of crowds and converge with significantly less observation time.
@@ 2 Problem Formulation
$s_{b}(t)=\alpha_{\text{s}}e^{j\theta_{s}}+\sum_{m\in\Omega}\alpha_{m}e^{j\left(\frac{2\pi}{\lambda}\psi_{m}\int v_{m}(t)dt+\frac{2\pi}{\lambda}d_{m}\right)},$ (1)
@@
$p(t)=\sum_{m\in\Omega}A_{m}\cos\left(\frac{2\pi\psi_{m}}{\lambda}\int v_{m}(t)dt+\Delta\mu_{m}\right),\vskip-3.0pt$ (2)
@@
Figure 2: Sample WiFi experiment from a stationary crowd of three individuals to demonstrate the difference between binary and non-binary bandwidth-based motion representations – (a) spectrogram of the received WiFi signal capturing real-time fidget activities, (b) binary fidget representation obtained via thresholding, which provides a coarse, on/off encoding of motion, (c) proposed bandwidth-based representation enabling a more fine-grained and informative motion description, if it could be related to the crowd size in a principled way. In this paper, we show how to methodically extract the crowd size from the bandwidth.
@@ 4.1 Relating Body Part Speeds of an Individual to Bandwidth
4.2 Modeling the Observed Bandwidth of a Stationary Crowd of Size NN
@@
We begin by expressing the cumulative distribution function (CDF) for a single individual as: $F_{BW}(x)=P(BW\leq x)=\int_{0}^{x}f_{BW}(\xi)\,d\xi$. In order to extend this model to a crowd of $N$ individuals, we apply the following transformation. We assume that the bandwidth contributions from each individual are independent and identically distributed (i.i.d.), which is a reasonable assumption as individuals do not typically fidget in a correlated manner. Then, at each point in time, the bandwidth contribution from each person follows the distribution $f_{BW}(x)$, with the observed final bandwidth for a crowd determined by the maximum among the $N$ independent contributions. This follows from the fact that signal components from different individuals combine linearly, and the dominant bandwidth is dictated by the fastest-moving body part across the group. Specifically, for i.i.d. random variables, we have the following for the cumulative bandwidth distribution for a crowd of $N$ individuals:
@@
$=\prod_{i=1}^{N}P\left(BW_{i}\leq x\right)\quad\text{(}BW_{i}\text{ are independent)}$
$=\left(P\left(BW\leq x\right)\right)^{N}\mkern 1.0mu\text{(}BW_{i}\text{ are identically dist.)}$
@@ 5.1 Experiment Settings
We conducted our experiments in two distinct indoor environments: a conference room and an office space, as shown in Fig. 1. We chose indoor spaces, as they present considerably more challenges due to the higher level of multipath, as compared to outdoor spaces. Moreover, the two indoor spaces are further chosen to represent different degrees of multipath. The conference room ($8$ m x $10.5$ m) is moderately furnished and exhibits a smaller level of multipath propagation, thus offering a relatively clean RF environment. In contrast, the office ($3$ m x $4$ m) is densely populated with metallic furniture, shelves, and electronics, resulting in a highly cluttered environment with significant multipath effects. This diversity allows us to evaluate the robustness of our approach in both relatively challenging and considerably challenging RF conditions.
In each environment, we conducted a series of experiments with varying crowd sizes and seating arrangements. Specifically, a subset of 13 participants, including 11 males and 2 females, is recruited and asked to be seated for reading. The configuration of the seating, including the number of rows of chairs, spacing between individuals, and alignment with respect to the TX-RX link, is varied across trials to ensure generalization across spatial layouts. The participant distance from the TX-RX link varied between 1.5 m and 4.5 m, and the minimum seat spacing between participants ranged from 0.5 m to 1 m across all experiments. Each experiment lasted around 3 minutes. We note that the small variation around experiment times is due to experimental conditions, external interventions, and participant preparations. The experiments were held across multiple days.
@@ 6.2 Simulation-Based Results to Test Larger Crowd Sizes
So far, we have shown real experimental results with crowds up to and including 13 individuals. To showcase our pipeline’s performance with even larger crowds, we next present a novel framework to efficiently simulate and test with larger crowd sizes. This pipeline synthesizes crowd bandwidths directly from readily available videos. Leveraging this video-to-bandwidth modeling eliminates the need for a full-body crowd wave simulator (such as those used in [46]), which is difficult to make fully realistic in terms of wave interactions.
More specifically, to simulate synthetic bandwidth observations for larger crowd sizes, we use the video-to-bandwidth pipeline of Section IV-C and public videos of single seated individuals. Each video is segmented into 3-minute non-overlapping segments. Each segment is then processed through the video-tobandwidth pipeline, which yields the corresponding bandwidth measurements for the seated individual as if a transceiver were present. To synthetically generate test data for any crowd size $N$, we then randomly sample $N$ of these segments and combine them by taking the maximum bandwidth value across all $N$ segments at each time instance. This approach efficiently models the observed bandwidth expected from a group of $N$ independently fidgeting, seated individuals, where the overall observed bandwidth is dictated by the highest contributor at any given time. This aggregation mimics the dominant bandwidth behavior of real multi-person environments. We repeat this sampling strategy for $N\in{2,3,...,20}$ with as many synthetic test samples generated per $N$ as the number of unique combinations that can be created from the video pool. Fig. 4 summarizes our proposed method for synthetic test data generation.
@@
In order to mitigate the bias towards any specific random selection, we repeat the entire cross-validation process five times with different randomizations and report the average results. In total, for each crowd size, we evaluate our pipeline on $296$ unique 3-minute segments per crowd size, yielding a total test set size of $20\times 296=5920$ empirical test instances. We note that each data segment is unique. However, during the extrapolation to larger crowd sizes, different segments belonging to the same video may be randomly picked as part of the same synthetic crowd data sample.
Table 4 reports the MAE for each crowd size. Our system achieves an overall NMSE of $0.054$ and an overall MAE of $1.47$ over the simulated dataset. These results further showcase the potential of our bandwidth-based approach with even larger crowd sizes.
@@ 6.3 Computation Time
7 Discussion and Future Directions
We next discuss several aspects of our proposed work and possible future directions.
7.1 Impact of $\psi$
As mentioned in Sec. 5.3, the TX-RX separation is intentionally chosen to be small (0.5 m) in our experiments. This design choice ensures that $\psi$ remains very close to its upper bound (2) for all realistic crowd distances. For example, when a person is as close as 1.5 m to the link, $\psi=1.97$. Such a close TX-RX setup and the corresponding $\psi\approx 2$ is commonly used in the RF sensing literature. The high performance of our pipeline further confirms the validity of this assumption in our setup. If the TX and RX are not close to each other, $\psi$ can deviate from 2. RF sensing under such a setup is less studied and constitutes an important avenue for future work.
7.2 Independence of the received bandwidth across individuals
In Sec. 4.2, we took the fidgeting patterns of different individuals in a seated crowd to be uncorrelated. To validate this assumption, we conducted two experiments in which six participants were video recorded while reading. Using the MediaPipe pose extractor together with the proposed video-to-bandwidth pipeline (Sec. 4), we generated individual bandwidth time series and computed pairwise Pearson correlation coefficients across participants. The average correlation coefficient magnitude was $0.142$ in the first experiment and $0.118$ in the second one. The resulting low correlation values further confirm that individual fidgeting behaviors are largely uncorrelated across participants, supporting the i.i.d. assumption of Sec. 4.2.
7.3 Impact of the anomaly detector threshold
In Sec. 5.4, we set the anomaly detection rule such that any window exhibiting a reconstruction error greater than 50% above the average training reconstruction error is flagged as anomalous and excluded from further analysis. In this section, we conduct a sensitivity analysis by varying the threshold across a range of values. Figure 5 shows the resulting MAE of our pipeline as the anomaly detection threshold is swept from 1.0 to 2.0 times the training reconstruction error. As can be seen, our pipeline’s performance is robust with respect to the threshold, with the MAE remaining within the narrow range of 0.97–1.04 when the anomaly detection threshold lies between 1.25 and 1.75 times the training reconstruction error.
Figure 5: Sensitivity of the proposed pipeline to the anomaly detection threshold. The threshold is expressed as a multiple of the average training reconstruction error (e.g., 1.50 corresponds to 50% above the training reconstruction error). MAE remains stable over a wide range of thresholds, demonstrating the robustness of the proposed pipeline.
7.4 Applicability Beyond Reading
While our experiments focus on cases where the crowd is engaged in reading, the proposed pipeline does not rely on reading as the activity. Instead, we have proposed a method to exploit the fidget statistics of a seated crowd to infer its size. Therefore, the proposed framework can generalize to other seated activities (e.g., studying, typing, or watching a presentation) by utilizing the relevant activity videos.